Determine the critical values for the confidence interval for the population variance from the given values. Round your answers to three decimal places. n=11 and α=0.1.
Answer #1

df = n -1 = 11 - 1 = 10

Chi square critical values at 0.1 significance level with 10 df = 3.940 , 18.307